OpenStudy (anonymous):

solve for x. x+2/3=-4 @Liv1234

OpenStudy (liv1234):

Alright, are there any options?

OpenStudy (anonymous):

A. –3 B. 1 C. –10 D. –14

OpenStudy (liv1234):

okay, so do you know how to solve these?

OpenStudy (anonymous):

sorta

OpenStudy (liv1234):

So, what you wanna do is plug in each number to the X value and see what each one gets you. Let's start with -3.

OpenStudy (anonymous):

okay

OpenStudy (liv1234):

\[-3+2/3=?\]

OpenStudy (anonymous):

so -3+2/3=4

OpenStudy (liv1234):

Correct, what do you get after plugging the -3 in? Hint: it's not -4.

OpenStudy (liv1234):

\[-2.3\] (the three is repeating)

OpenStudy (anonymous):

lol um frist add then divide

OpenStudy (anonymous):

try to see if it makes -4

OpenStudy (liv1234):

Oh my, wait, we have to divide first and then add, my mistake.

OpenStudy (anonymous):

okay

OpenStudy (liv1234):

I got the same thing for A when I divided first and then added, so it's not A

OpenStudy (anonymous):

okay

OpenStudy (liv1234):

\[1+2/3\]

OpenStudy (anonymous):

that equals 1 right?

OpenStudy (liv1234):

Sorry, I was solving haha.

OpenStudy (anonymous):

@fishejac000

OpenStudy (liv1234):

So, 1+2=3

OpenStudy (liv1234):

3/3=1

OpenStudy (liv1234):

So, it's not A or B.

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

i put accidently put this guy in this sorrt ddint want to do that

OpenStudy (anonymous):

okay so its not b

OpenStudy (anonymous):

c?

OpenStudy (liv1234):

Correct and it's not C because it doesn't equal -4, so what does that leave you with?

OpenStudy (anonymous):

d!!!!!

OpenStudy (liv1234):

\[-14+2=-12/3=\]

OpenStudy (liv1234):

Correct!(: Medal?

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

i have another one

OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16

OpenStudy (anonymous):

A. 21 B. 3 C. 36 D. –3

OpenStudy (liv1234):

Hint: PEMDAS

OpenStudy (anonymous):

real quick is it a

OpenStudy (liv1234):

Okay, so 16/2=8*4=32-10/5+6

OpenStudy (liv1234):

Hint: It's not A or D

OpenStudy (anonymous):

so let me calculate this

OpenStudy (liv1234):

It's 36 (:

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

its -14/3